# Normative feeling: socially patterned affective mechanisms

### v3

@@ 4.3 Mood mechanism that emerges with punishment avoids population collapse
In the previous section we showed that punishment patterned the mood mechanism, resulting in a behaviour that is socially rational - i.e. eat less when losing energy and eat more when gaining energy. This means the agents avoid eliciting punishment from each other. Now we will examine what this does on the group level to the population dynamics and ascertain whether this is indeed beneficial to the group and therefore prosocial.
In the population plots for the competition condition Figure (Figure 6a), we see that populations grow when there is plenty of resource, however, during periods of low resource the populations collapse due to overexploitation of the resource during the growth phase.
In contrast, for a subset of runs in the punishment condition agent populations do not overgrow during periods of high resource and therefore do not experience the same collapse of population due to overexploitation Figure 6b. This is due to agents punishing each other, taking energy away from each other, and therefore not allowing the overgrowth that results in collapse of the population later on. We argue that this is a prosocial outcome form the perspective that overgrowth and then collapse results in many agent deaths. This being said, an issue that emerges: Since punishment is so costly to the population(Figure 6e), it results in an extreme inefficiency that means the populations remain very small.
@@
Figure 6: This figure depicts the average population level traits for different simulation runs over time. Comparing the competition and punishment condition. N.B. In the punishment condition we plotted only runs that evolved high punishment and conserved energy compared to the competition condition. a) Population of runs in competition condition b) Population of runs in punishment condition c) Hunger of runs in competition condition d) Hunger of runs in punishment condition coloured lined depicts average of trait for individual runs and black line is average over all runs e) Injuries from sanction in punishment condition: this is the average energy lost to punishment across the population.
@@ 4.4 Mood injection to causally confirm effect of mood on behaviour
To do this, we artificially "injected" positive mood into the agents to see if there was a net effect on the behaviour. By this we mean artificially making the mood of every agent positive to see how agents change behaviour when they are in a positive mood. We did this every 500 time steps, "injecting" 200 units of positive mood for 200 time steps into each agent and these graphs show how this affected behaviour Figure 7. We confirm that agents in the punishment condition increase their $\mu_{eat}$ when they are in a good mood and decrease it when they are in low mood, and that they punish more in a good mood and punish less in a bad mood. We see that, especially in the $\mu_{sanction}$ graph that not all populations respond in the same way when positive mood mechanism is injected e.g. some decrease their amount of punishment. However, we can see that on average that injecting mood mechanism has the effect described above (eat more and punish more in a good mood, eat less and punish less in a bad mood) Figure 7. Therefore, we can conclude that punishment does indeed cause our model of mood to become patterned as seen here. This population level connection between mood and behaviour opens up the possibility of the agents influencing each other with mood signals. This is since now, agents in the population now have a similar behavioural response to one another when they are in similar mood. For example, if an agent could emit a signal that made others feel "bad" by lowering their mood, they would get them to consume less resource and vice versa. This would be a more sophisticated method to regulate other agents behaviour then physically punishing them. Signalling of this type is a more sophisticated way that organisms regulating each others behaviour, and avoids the inefficiencies and detrimental aspects of physical punishment [5]. Therefore punishment (axiomatic normativity), by patterning mood mechanism on the population level, has set the stage for the evolution of higher order normative signalling.
Figure 7: Mood injection: the average value for each population is plotted for two traits: $\mu_{eat}$ above and $\mu_{sanction}$ below. The average across the population is plotted as a black line. Injections of positive mood occur every 500 time steps for 200 steps. Since positive mood (feeling good) is defined by whether the weights for avg-$W_{energy-gain}$(average energy gain) are positive or negative, we injected negative mood if it was negative (to create a net positive) and positive if it was positive (again for net positive). This means that for any case we are effectively injecting positive mood into the system
@@ 5 Discussion
In this article asked how our moods are entangled with social norms i.e. how did social preferences emerge in our moods evolutionarily? We argue, that since norm like processes (minimal normativity) are replete across phyla, that it could be a possible selection pressure that lead to social preferences evolving in mood mechanisms, despite this many accounts of mood remain individualist.
We decided to address this with an agent based model in which agents have a mood mechanism whose vital parameters could be evolved, therefore allowing us to study the effect of different aspects of the physical and social environment (specifically axiomatically normative punishment) affects evolution of mood mechanisms in an artificial setting.
@@ 5.1 Summary of results
  • 1b. We verify that punishment (explicit regulation of another’s cultural or genetic phenotypic fitness), creates multiple equilibria and is therefore qualitatively different from other forms of social maintenance/social behaviour e.g. competition, and therefore fundamentally normative.
  • 2. Use our axiomatic normativity definition to model the effect of normative punishment on the evolution of mood. We show that this could explain the emergence of social preference i.e. how can mood systems become patterned in such a way as to be altruistic (not just the binary uptake of a variable or not, as it’s usually modelled).
  • 3. Show that minimally normative processes imprint themselves on the mood mechanism and allow mood mechanisms to become normative themselves, setting the stage for evolution of mood signalling mechanisms.
@@ 5.2 Discussion of results
We show that considering normative punishment in the evolution of mood has a strong effect, producing a qualitatively different mood mechanism compared to considering just competition. We distinguish between competition and punishment, showing that the latter is axiomatically normative. We define an axiomatically normative process as a processes that gets agents to converge on a behaviour while producing multiple behavioural equilibria instead of just one. Thus capturing a key aspect of normativity; a certain degree of arbitrariness. The fact that normative punishment is widely conserved across phyla [36], even in simple organisms like insects [46] and bacteria [24], should urge us to consider the effects of axiomatically normative processes on the evolution even in ancient mechanisms such as mood and other cognitive processes.
Using an evolvable model of mood we show that evolution and punishment alone can result in the encoding of a social preference in a mood mechanism. This mood mechanism takes into account the negative social consequences of taking up too many resources, encoded in a distributed manner across the weights of the mood mechanism, and gets the agent to act in a socially adaptive manner. This is in contrast to other models of social preference, where preferences are often hardcoded into models (e.g normative concern/ guilt [1, 11] and explanations about how they emerge evolutionary from purely material concerns i.e. purely from individual level self interest are thin on the ground and assume trivial mappings between genes and behaviour that omit mechanism [10, 18]. We provide a possible mechanism for how social motivations/preferences can evolve from purely material incentives and normative social interaction. In other words, social preferences to are co-constructed in a decentralised manner through agents manipulating each others fitness through normative punishment. This also goes to show that there is no objectively good mood mechanism but that it depends on the interactions between the agent and the environment (in this case social environment). In other words, there is no "right way" to feel, it is subjective, it depends on what is adaptive given the context.
@@
The mood mechanism allows multiple concerns to be encoded in one i.e. mood reflecting both lack of food but also level of punishment. One could start thinking of mood as funnelling certain goals into one. One may feel bad when they do not get what they are individually interested in but also when socially shamed (i.e. ones mood depends on what others think of you). If an agent maximises for feeling "good" then they are implicitly capturing multiple goals in one "utility" to maximise. This can be the basis for brains or in this case simple mood mechanisms becoming socially entangled and therefore socially rational [22]. Furthermore, agents withdrawing from taking from the shared resource when in a "bad" mood, in order to avoid punishment, can be interpreted as a depressive mood. According to the social risk hypothesis, depression represents an adaptive response to the perceived threat of exclusion from important social relationships that, over the course of evolution, have been critical to maintaining an individual’s fitness prospects [3, 2].
We show through causal probing of our simulation. That invoking negative or positive mood in the evolved agents results in agents reacting in a predictable manner, eating less when they are in a bad mood and more when they are in a positive mood ". This opens up signalling to each other to change their behaviour with the threat of punishment [5] i.e. direct effect on mood through negative signals (yelling at someone). This would mean that you could get the desired population regulation effect without the physical damage from the sanctions Figure 7. This is possible because including punishment in the first place has resulted in social entanglement. That is, alignment of the way mood mechanisms work meaning agents are socially entangled and have a common emotional response. This means that they can in turn use emotional (positive and negative) signals to influence each other. In effect, punishment has imprinted a kind of "normative logic" into the mood mechanism that allows more sophisticated mind shaping to take place without the need for agents to physically punish each other, which is costly. This kind of mind shaping is commonly throughout to occur in organisms only with complex cognitive faculties like Theory of Mind [17, 49] but we suggest a plausible path were it can emerge from the simple ingredients of mood and normative punishment .
@@ 5.3 Advantage of our model vs other models
Our model does not use discrete motivations as assumed in other models i.e. [1, 29] but allows them to be encoded in a distributed manner across a network. On the other hand, different from traditional neural network models which, while more abstract and with fewer assumptions, are harder to relate to the real world due to their opaqueness. Our model can relate back to real motivations through the use of affective language while still seeing how social preferences are “smeared” across a network instead of presupposing discrete motivations. Therefore our model is a nice middle ground between these approaches.
@@ 5.4 Limitations
We have only included punishment (negative explicit regulation of others’ fitness in our model. But, we also know that positive regulation may exist through incensation of good behaviour for example, whose inclusion may allow for richer dynamics to emerge. Moreover, we have a single dimensional view of mood and there exist different models that have at least two dimensions other than valance and magnitude [33, 43]. In the future, we can expand our model to see the effects of minimal normativity on these models of mood.
@@ 5.5 Future work
We showed that that punishment and mood in the preserve of evolution resulted in alignment of mood mechanisms in a population. Meaning agents are socially entangled and have a common emotional response which allows for higher order mindshaping [49] through signalling [5]. That is, direct effect on mood through negative signals (yelling at someone) to regulate one another. Future work could use simulations to see if it is possible to evolve mood signalling. This would help answer if indeed this sophisticated mind shaping behaviour can emerge from the simple ingredients of normativity and mood. Possible complications of having evolved mood signalling would be interesting to consider. For example it opens up the possibility of emotional free-riders i.e. agents that manage to "unlearn" or "unevolve" their response to the punishment signal to regulate themselves and start acting selfishly again. Or you may even get others abusing the signal for their own gain, reducing others greed in order to reek the benefits, a sort of mood manipulation. This may impede mood signalling evolving at all, with agents evolving to have completely unpredictable responses in order to not be manipulated. For evolution to overcome this roadblock we may need to posit e.g. honest signalling mechanisms or commitment devices for something like this to evolve [32, 27].

### v4

@@ 4.3 Mood mechanism that emerges with punishment avoids population collapse
In the previous section, we showed that punishment patterned the mood mechanism, resulting in a behaviour that is socially rational - i.e. eat less when losing energy and eat more when gaining energy. This means the agents avoid eliciting punishment from each other. Now we will examine what this does on the group level to the population dynamics and ascertain whether this is indeed beneficial to the group and therefore prosocial.
In the population plots for the competition condition (Figure 6a), we see that populations grow when there is plenty of resource; however, during periods of low resource the populations collapse due to overexploitation of the resource during the growth phase.
In contrast, for a subset of runs in the punishment condition, agent populations do not overgrow during periods of high resource and therefore do not experience the same collapse of population due to overexploitation, Figure 6b. This is due to agents punishing each other, taking energy away from each other, and therefore not allowing the overgrowth that results in collapse of the population later on. We argue that this is a prosocial outcome from the perspective that overgrowth and then collapse result in many agent deaths. This being said, an issue that emerges: Since punishment is so costly to the population(Figure 6e), it results in an extreme inefficiency that means the populations remain very small.
@@
Figure 6: This figure depicts the average population level traits for different simulation runs over time. Comparing the competition and punishment conditions. N.B. In the punishment condition, we plotted only runs that evolved high punishment and conserved energy compared to the competition condition see SI for full runs). a) Population of runs in competition condition b) Population of runs in punishment condition c) Hunger of runs in competition condition d) Hunger of runs in punishment condition. The coloured line depicts the average of trait for individual runs and black line is average over all runs e) Injuries from sanction in punishment condition: this is the average energy lost to punishment across the population.
@@ 4.4 Mood injection to causally confirm effect of mood on behaviour
To do this, we artificially "injected" positive mood into the agents to see if there was a net effect on the behaviour. By this, we mean artificially making the mood of every agent positive to see how agents change their behaviour when they are in a positive mood. We did this every 500 time steps, "injecting" 200 units of positive mood for 200 time steps into each agent, and these graphs show how this affected behaviour Figure 7. We confirm that agents in the punishment condition increase their $\mu_{eat}$ when they are in a good mood and decrease it when they are in a low mood, and that they punish more in a good mood and punish less in a bad mood. We see that, especially in the $\mu_{sanction}$ graph, not all populations respond in the same way when a positive mood is injected, e.g. some decrease their amount of punishment. However, we can see that on average, injecting mood mechanism has the effect described above (eat more and punish more in a good mood, eat less and punish less in a bad mood), Figure 7. Therefore, we can conclude that punishment does indeed cause our model of mood to become patterned as seen here. This population-level connection between mood and behaviour opens up the possibility of the agents influencing each other with mood signals. This is because agents in the population now have a similar behavioural response to one another when they are in a similar mood. For example, if an agent could emit a signal that made others feel "bad" by lowering their mood, they would get them to consume less resource and vice versa. This would be a more sophisticated method to regulate other agents’ behaviour than physically punishing them. Signalling of this type is a more sophisticated way that organisms regulate each other’s behaviour, and avoids the inefficiencies and detrimental aspects of physical punishment [5]. Therefore, punishment (axiomatic normativity), by patterning mood mechanism on the population level, has set the stage for the evolution of higher-order normative signalling.
Figure 7: Mood injection: the average value for each population is plotted for two traits: $\mu_{eat}$ above and $\mu_{sanction}$ below. The average across the population is plotted as a black line. Injections of positive mood occur every 500 time steps for 200 steps. Since positive mood (feeling good) is defined by whether the weights for avg-$W_{energy-gain}$(average energy gain) are positive or negative, we injected negative mood if it was negative (to create a net positive) and positive if it was positive (again for net positive). This means that for any case, we are effectively injecting positive mood into the system/
@@ 5 Discussion
In this article asked how our moods are entangled with social norms, i.e. how did social preferences emerge in our mood mechanisms evolutionarily? We argue that since norm-like processes (minimal normativity) are replete across phyla, it could be a possible selection pressure that leads to social preferences evolving in mood mechanisms; despite this, many accounts of mood remain individualist.
We decided to address this with an agent-based model in which agents have a mood mechanism whose vital parameters could be evolved, therefore allowing us to study the effect of different aspects of the physical and social environment (specifically, axiomatically normative punishment) on the evolution of mood mechanisms in an artificial setting.
@@ 5.1 Summary of results
  • 1b. We verify that punishment (explicit regulation of another’s cultural or genetic phenotypic fitness) creates multiple equilibria and is therefore qualitatively different from other forms of social maintenance/social behaviour, e.g. competition, and therefore fundamentally normative.
  • 2. Use our axiomatic normativity definition to model the effect of normative punishment on the evolution of mood. We show that this could explain the emergence of social preference, i.e. how mood systems become patterned in such a way as to be altruistic (not just the binary uptake of a variable or not, as it’s usually modelled).
  • 3. Show that minimally normative processes imprint themselves on the mood mechanism and allow mood mechanisms to become normative themselves, setting the stage for the evolution of mood signalling mechanisms.
@@ 5.2 Discussion of results
We show that considering normative punishment in the evolution of mood has a strong effect, producing a qualitatively different mood mechanism compared to considering just competition. We distinguish between competition and punishment, showing that the latter is axiomatically normative. We define an axiomatically normative process as a process that gets agents to converge on a behaviour while producing multiple behavioural equilibria instead of just one. Thus, capturing a key aspect of normativity, a certain degree of arbitrariness. The fact that normative punishment is widely conserved across phyla [36], even in simple organisms like insects [46] and bacteria [24], should urge us to consider the effects of axiomatically normative processes on the evolution, even in ancient mechanisms such as mood and other cognitive processes.
Using an evolvable model of mood, we show that evolution and punishment alone can result in the encoding of a social preference in a mood mechanism. This mood mechanism takes into account the negative social consequences of taking up too many resources, encoded in a distributed manner across the weights of the mood mechanism, and gets the agent to act in a socially adaptive manner. This is in contrast to other models of social preference, where preferences are often hardcoded into models (e.g normative concern/ guilt [1, 11] and explanations about how they emerge evolutionary from purely material concerns i.e. purely from individual level self interest are thin on the ground and assume trivial mappings between genes and behaviour that omit mechanism [10, 18]. We provide a possible mechanism for how social motivations/preferences can evolve from purely material incentives and normative social interaction. In other words, social preferences are co-constructed in a decentralised manner through agents manipulating each other’s fitness through normative punishment. This also goes to show that there is no objectively good mood mechanism, but that it depends on the interactions between the agent and the environment (in this case social environment). In other words, there is no "right way" to feel; it is subjective, it depends on what is adaptive given the context.
@@
The mood mechanism allows multiple concerns to be encoded in one, i.e. mood reflecting both lack of food but also the level of punishment. One could start thinking of mood as funnelling certain goals into one. One may feel bad when they do not get what they are individually interested in, but also when socially shamed (i.e. ones mood depends on what others think of you). If an agent maximises for feeling "good", then they are implicitly capturing multiple goals in one "utility" to maximise. This can be the basis for brains, or in this case, simple mood mechanisms becoming socially entangled and therefore socially rational [22]. Furthermore, agents withdrawing from taking from the shared resource when in a "bad" mood, in order to avoid punishment, can be interpreted as a depressive mood. According to the social risk hypothesis, depression represents an adaptive response to the perceived threat of exclusion from important social relationships that, over the course of evolution, have been critical to maintaining an individual’s fitness prospects [3, 2].
We show through causal probing of our simulation that invoking negative or positive mood in the evolved agents results in agents reacting in a predictable manner: eating less when they are in a bad mood and more when they are in a positive mood ". This opens up signalling to each other to change their behaviour with the threat of punishment [5], i.e. direct effect on mood through negative signals (yelling at someone). This would mean that you could get the desired population regulation effect without the physical damage from the sanctions, Figure 7. This is possible because including punishment in the first place has resulted in social entanglement. That is, alignment of the way mood mechanisms work, meaning agents are socially entangled and have a common emotional response. This means that they can, in turn, use emotional (positive and negative) signals to influence each other. In effect, punishment has imprinted a kind of "normative logic" into the mood mechanism that allows more sophisticated mind shaping to take place without the need for agents to physically punish each other, which is costly. This kind of mind shaping is commonly thought to occur in organisms only with complex cognitive faculties like Theory of Mind [17, 49], but we suggest a plausible path where it can emerge from the simple ingredients of mood and normative punishment.
@@ 5.3 Advantage of our model vs other models
Our model does not use discrete motivations as assumed in other models, i.e. [1, 29], but allows them to be encoded in a distributed manner across a network. On the other hand, different from traditional neural network model,s which, while more abstract and with fewer assumptions, are harder to relate to the real world due to their opaqueness. Our model can relate back to real motivations through the use of affective language while still seeing how social preferences are “smeared” across a network instead of presupposing discrete motivations. Therefore, our model is a nice middle ground between these approaches.
@@ 5.4 Limitations
We have only included punishment (negative explicit regulation of others’ fitness in our model. But, we also know that positive regulation may exist through incensation of good behaviour, for example, whose inclusion may allow for richer dynamics to emerge. Moreover, we have a single-dimensional view of mood, and there exist different models that have at least two dimensions other than valence and magnitude [33, 43]. In the future, we can expand our model to see the effects of minimal normativity on these models of mood.
@@ 5.5 Future work
We showed that punishment and mood in the preserve of evolution resulted in alignment of mood mechanisms in a population. Meaning agents are socially entangled and have a common emotional response, which allows for higher-order mindshaping [49] through signalling [5]. That is, direct effect on mood through negative signals (yelling at someone) to regulate one another. Future work could use simulations to see if it is possible to evolve mood signalling. This would help answer whether indeed this sophisticated mind shaping behaviour can emerge from the simple ingredients of normativity and mood. Possible complications of having evolved mood signalling would be interesting to consider. For example, it opens up the possibility of emotional free-riders, i.e. agents that manage to "unevolve" their response to the punishment signal to regulate themselves and start acting selfishly again. Or you may even get others abusing the signal for their own gain, reducing others’ greed in order to reap the benefits, a sort of mood manipulation. This may impede mood signalling evolving at all, with agents evolving to have completely unpredictable responses in order not to be manipulated. For evolution to overcome this roadblock, we may need to posit, e.g. honest signalling mechanisms or commitment devices for something like this to evolve [32, 27].